# Theory-Grounded Evaluation of Human-Like Fallacy Patterns in LLM Reasoning

### v2

@@ 4 Experimental Results
Proportion of Errors that are Fallacies
@@
where $P$ is the set of all reasoning problems in our evaluation set. This measure represents the degree to which a model’s errors align with predictable human reasoning patterns rather than other forms of logical error. We find as model strength increases, so does the proportion of logically incorrect answers that are fallacies (Figure 1).
@@
Overall Correctness and other measures
Surprisingly, model strength as measured by LLM Elo score does not appear to have any correlation $(r=0.004,p=0.981)$, $(\rho=-0.04,p=0.777)$ with the model’s ability to produce correct answers on our dataset: see Table 3. Furthermore, exploratory analysis did not reveal any other statistically significant evidence of model Elo correlating or enjoying a monotonic ordinal relationship with absolute proportion of fallacies produced, whether exactly predicted by ETR or taking logical equivalences.
@@
Intervention study: Reversing Premise Order
While in classical logic the order in which premises are presented has no effect on normatively correct logical conclusions, an order-effect is commonly observed in human reasoners [39], where presenting premises in different orders elicits logically different responses. In addition to being a proxy measure for the non-classicality of human reasoners, this non-commutativity effect often blocks fallacies. We observe strong evidence that LLMs are similar in this regard: presenting the same questions with reversed premise order significantly blocks fallacy production. Table 4 reports a significant fallacy-blocking effect from reversing premises on a majority of models.
Table 4: 38 models selected for testing, listed by Chatbot Arena Score (general). The right-most column reports the percentage of fallacies per-model blocked (answer becomes logically correct) by reversing premise order, along with the results of a two-proportion $z$-test for each model to gauge the significance of the blocking effect.
@@
5 Conclusion and Discussion
Interpretation of Key Findings Our results reveal a statistically significant correlational trend in error composition: as language models advance in capability (proxied by Chatbot Arena Elo), a larger share of logically incorrect answers align with human errors predicted by ETR. While stronger language models demonstrate improved capabilities across many benchmarks, we do not observe increased overall correctness in this dataset. Moreover, we demonstrate order-effects for LLM reasoning analogous to those observed in humans. This indicates overlap between LLM error patterns and established human fallacy patterns on these tasks.
Methodological Considerations and Limitations While traditional ablation studies might seem appropriate for investigating this phenomenon, they face significant limitations in the current landscape of language models: the diversity of modern architectures, training approaches such as distillation  [40, 41, 42], and reasoning capabilities with test-time scaling  [43, 44, 45] means that simple parameters like model size may no longer serve as reliable proxies for model capability. Our use of Chatbot Arena scores provides a performance-based metric that naturally incorporates a diverse range of models, reflecting real-world ability rather than architectural specifics, similar to Ruan et al. [46].
@@ 5 Conclusion and Discussion
While our results demonstrate correlation between model capability and human-like errors, we acknowledge limitations in attributing causality. Alternative explanations include the possibility that stronger models are increasingly trained on human-produced reasoning traces that themselves contain these fallacies. The absence of improved logical accuracy with model capability may reflect ceiling effects in our test set or fundamental limitations in current training paradigms.
Theoretical Implications These findings challenge the assumption that scaling alone leads to more normatively correct reasoning systems. Instead, within our evaluation setting we observe greater overlap between model error patterns and characteristic human fallacies. This does not by itself imply shared internal processes; it highlights an evaluation axis (error composition) complementary to overall accuracy.
Our results provide a foundation for several research directions. For benchmark development, our approach offers a systematic method for generating reasoning problems where human-like fallacies are predicted, creating evaluation sets that could complement existing reasoning benchmarks. For capabilities, the identified fallacy patterns could serve as targeted training examples to fortify models against human-like reasoning errors while maintaining their general capabilities. As a tool for assessing engineering choices, our methodology provides a quantitative framework for evaluating the effectiveness of various interventions aimed at improving model reasoning, from prompt engineering to architectural modifications.
Broader Implications This observed trade-off in error composition has implications for AI alignment. As models become more capable, anticipating and mitigating human-like reasoning errors becomes increasingly important, especially in high-stakes contexts requiring reliable reasoning, such as medical diagnosis, legal analysis, or decision support [8, 9].
@@
Acknowledgments and Disclosure of Funding
The authors wish to thank the Cosmos Institute for their continued support.
@@ Appendix A A formal presentation of the Erotetic Theory
ETR is formulated in [10] in a set-theoretic language, and we reproduce the full definitions here, with the explicit permission of the authors. We remark that Definitions 1 and 11 contain the parameters of the theory, i.e. the basic atomic propositions, the constants and function symbols, and a set logical and extralogical axioms. The central definition is that of a View, Definition 9. Following that are definitions of the basic operations of ETR, in particular Update (Definition 23) and Query (Definition 30) whose use in PyETR were demonstrated in Section 2.
@@
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{A}=(\arg\max\limits_{\gamma\in\Gamma}\ \Delta[\{\{p\}:p\in\gamma\}]^{\mathcal{A}P})^{\Theta}_{[R][I]}$

### v3

@@ 4 Experimental Results
4.1 Proportion of Errors that are Fallacies
@@
where $P$ is the set of all reasoning problems in our evaluation set. This measure represents the degree to which a model’s errors align with predictable human reasoning patterns rather than other forms of logical error. We find as model strength increases, so does the proportion of logically incorrect answers that are fallacies. We observe this trend consistently across model Elos scores (Figure 1), as well as HELM composite benchmark scores and training compute estimates (Appendix C).
@@
4.2 Error Analysis
We manually inspected LLM responses from our experiment and found many patterns in line with ETR’s predictions. Models often fixate on repeated objects, ignoring disjunctive alternatives to conclude fallacious categorical answers. Models also instantiate existential quantifiers with objects from premises and generally mis-scope quantifiers when presented with repeat information. Appendix D shows some illustrative examples of model fallacy cases.
4.3 Overall Correctness and Other Measures
Surprisingly, model strength as measured by LLM Elo score does not appear to have any correlation $(r=0.004,p=0.981)$, $(\rho=-0.04,p=0.777)$ with the model’s ability to produce correct answers on our dataset (Table 3). Furthermore, exploratory analysis did not reveal any other statistically significant evidence of model Elo correlating or enjoying a monotonic ordinal relationship with absolute proportion of fallacies produced, whether exactly predicted by ETR or taking logical equivalences.
@@ 4.3 Overall Correctness and Other Measures
4.4 Intervention Study: Reversing Premise Order
While in classical logic the order in which premises are presented has no effect on normatively correct logical conclusions, an order-effect is commonly observed in human reasoners (Girotto et al., 1997), where presenting premises in different orders elicits logically different responses. In addition to being a proxy measure for the non-classicality of human reasoners, this non-commutativity effect often blocks fallacies. We observe strong evidence that LLMs are similar in this regard: presenting the same questions with reversed premise order significantly blocks fallacy production. Table 4 reports a significant fallacy-blocking effect from reversing premises on a majority of models. Appendix E shows an example of premise reversal blocking an ETR-predicted fallacy.
Table 4: 38 models selected for testing, ordered by Chatbot Arena Elo. The right-most column reports the percentage of fallacies per-model blocked (answer becomes logically correct) by reversing premise order, along with the results of a two-proportion $z$-test for each model to gauge the significance of the blocking effect.
@@ 4.4 Intervention Study: Reversing Premise Order
5 Discussion and Conclusion
5.1 Interpretation of Key Findings
Our results reveal a statistically significant correlational trend in error composition: as language models advance in capability (proxied by Chatbot Arena Elo), a larger share of logically incorrect answers align with human errors predicted by ETR. While stronger language models demonstrate improved capabilities across many benchmarks, we do not observe increased overall correctness in this dataset. Moreover, we demonstrate order-effects for LLM reasoning analogous to those observed in humans. This indicates overlap between LLM error patterns and established human fallacy patterns on these tasks.
5.2 Methodological Considerations and Limitations
While traditional ablation studies might seem appropriate for investigating this phenomenon, they face significant limitations in the current landscape of language models: the diversity of modern architectures, training approaches such as distillation (Gou et al., 2021; Gupta and Agrawal, 2022; Xu et al., 2024), and reasoning capabilities with test-time scaling (DeepSeek-AI, 2025; El-Kishky et al., 2024; Xu et al., 2025) mean that simple parameters like model size may no longer serve as reliable proxies for model capability. Our use of Chatbot Arena scores provides a performance-based metric that naturally incorporates a diverse range of models, reflecting real-world ability rather than architectural specifics, similar to the observational approach in Ruan et al. (2024).
@@ 5.2 Methodological Considerations and Limitations
While our results demonstrate correlation between model capability and human-like errors, we acknowledge limitations in attributing causality. Alternative explanations include the possibility that stronger models are increasingly trained on fallacious human-produced reasoning traces, or that post-training methods like RLHF cause convergence in reasoning behaviour. The absence of improved logical accuracy with model capability may reflect ceiling effects in our test set or fundamental limitations in current training paradigms.
5.3 Theoretical Implications
These findings challenge the assumption that scaling alone leads to more normatively correct reasoning systems. Instead, within our evaluation setting we observe greater overlap between model error patterns and characteristic human fallacies. This does not by itself imply shared internal processes; it highlights an evaluation axis (error composition) complementary to overall accuracy. We do not advance any particular mechanistic claim, as our results only support an empirical characterization.
Nonetheless, our results provide a foundation for several research directions. For benchmark development, our approach offers a systematic method for generating reasoning problems where human-like fallacies are predicted, creating evaluation sets that could complement existing reasoning benchmarks. For capabilities, the identified fallacy patterns could serve as targeted training examples to fortify models against human-like reasoning errors while maintaining their general capabilities. As a tool for assessing engineering choices, our methodology provides a quantitative framework for evaluating the effectiveness of various interventions aimed at improving model reasoning, from prompt engineering to architectural modifications. Follow-up work could investigate other predictive features, like reasoning trace length, answer diversity, and trends within model families.
5.4 Broader Implications
Our observed trade-off in error composition has implications for AI alignment. As models become more capable, anticipating and mitigating human-like reasoning errors becomes increasingly important, especially in high-stakes contexts requiring reliable reasoning, such as medical diagnosis, legal analysis, or decision support (Jacobs, 2021; Bommasani et al., 2022).
@@ 5.4 Broader Implications
Acknowledgements and Disclosure
Members of the Laboratory for Human-Centered AI are supported by funding from the Cosmos Institute. A.K.R. and R.O.K. were supported by a Cosmos Fellowship for the duration of this project. R.O.K. is additionally supported by the Clarendon Scholarship and the Jesus College Old Members’ Scholarship.
@@ Appendix A A formal presentation of the Erotetic Theory
ETR is formulated in Koralus (2022) in a set-theoretic language, and we reproduce the full definitions here, with the explicit permission of the authors. We remark that Definitions 1 and 11 contain the parameters of the theory, i.e. the basic atomic propositions, the constants and function symbols, and a set logical and extralogical axioms. The central definition is that of a View, Definition 9. Following that are definitions of the basic operations of ETR, in particular Update (Definition 23) and Query (Definition 30) whose use in PyETR were demonstrated in Section 2.
@@
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{A}=(\operatorname*{arg\,max}\limits_{\gamma\in\Gamma}\ \Delta[\{\{p\}:p\in\gamma\}]^{\mathcal{A}P})^{\Theta}_{[R][I]}$
@@
Appendix B List of thematic templates
Our 12 natural language themes are provided below. We provide the object and predicate names used when mapping to natural language with each theme, as well as the prompt used before the premises in each question.
We manually craft each list of objects and predicates to allow that arbitrary mappings remain semantically coherent. This requires all combinations of predicates to be commensurable, and for the objects to bear no relations to each other or to specific predicates. For example, in the “Cards” theme, the predicates “red,” “square,” and “marked” are commensurable (it is possible for a card to be red and not square, or red, square, and marked, and so on), and predicates with special relations to certain cards (like “is a face card” or “is even”) are excluded.
Cards
Prompt
I’m playing a card game against the computer. It’s an unusual game with an unusual deck of cards. Each card might be multiple colors or multiple shapes, for example. I have some clues about what’s going on, and I need to figure some more things out through logical reasoning. Here’s what I know so far:
Objects
  • the ace
  • the one
  • the two
  • the three
  • the four
  • the five
  • the six
  • the seven
  • the eight
  • the nine
  • the ten
  • the jack
  • the queen
  • the king
Predicates
  • red
  • square
  • marked
  • yellow
  • round
  • castable
Elements
Prompt
I’m working in a materials science lab and we’ve gotten some puzzling results. I need to use logical reasoning to figure out what’s going on. Here’s what I know so far:
Objects
  • elementium
  • zycron
  • phantasmite
  • mystarium
  • oblivium
  • luminite
  • darkonium
  • velocium
  • quasarium
  • chronium
  • aetherium
  • voidite
  • pyroflux
  • cryon
  • nebulium
  • solarium
  • eclipsium
  • stellarite
  • fluxium
  • gravitron
  • xylozine
  • ignisium
  • aurorium
  • shadowium
  • plasmor
  • terranite
  • harmonium
  • zenthium
  • celestium
  • radionite
Predicates
  • radioactive
  • luminescent
  • superconductive
  • magnetic
  • corrosive
  • volatile
  • plasma-like
  • gravity-enhancing
  • dimension-warping
  • time-dilating
  • self-repairing
  • shape-shifting
  • anti-matter reactive
  • bio-compatible
  • crystal-forming
  • acidic
  • alkaline
  • liquid at room temperature
  • gaseous under high pressure
  • solid in vacuum
  • quantum-stable
  • emotion-reactive
  • thermal-conductive
  • electrically insulating
  • transparent to visible light
  • dark energy absorbing
  • neutrino-emitting
  • anti-gravity generating
  • sound-absorbing
Planets
Prompt
I’m an astronomer studying newly discovered celestial bodies. I’ve made some observations and I need to use logical reasoning to figure out what’s going on. Here’s what I know so far:
Objects
  • planet X
  • planet Y
  • planet Z
  • asteroid A
  • asteroid B
  • comet 1
  • comet 2
  • moon 1
  • moon 2
  • moon 3
Predicates
  • rocky
  • gaseous
  • ringed
  • within a habitable zone
  • orbited by satellites
  • in retrograde orbit
  • elliptically-orbiting
  • visible to the naked eye
  • atmospheric
  • shielded by a magnetic field
  • tidally locked
Magical creatures
Prompt
I’m a magizoologist studying unusual creatures in my sanctuary. I need to understand their behaviors and characteristics through logical reasoning. Here’s what I’ve observed so far:
Objects
  • phoenixling
  • shadowdrake
  • moonwolf
  • crystalspider
  • stormgriffin
  • dreamweaver
  • frostwyrm
  • sunlion
  • etherealsnake
  • timefox
Predicates
  • firebreathing
  • shadow-walking
  • moonlight-glowing
  • crystal-forming
  • storm-controlling
  • dream-affecting
  • ice-generating
  • light-emitting
  • phase-shifting
  • time-bending
  • telepathic
  • aura-healing
  • able to turn invisible
  • shapeshifting
Enchanted artifacts
Prompt
I’m an arcane researcher cataloging mysterious magical items. I need to understand their properties through careful logical analysis. Here’s what I’ve documented so far:
Objects
  • Timekeeper’s Compass
  • Void Mirror
  • Dreamcatcher Ring
  • Starlight Pendant
  • Shadow Cloak
  • Crystal Orb
  • Phoenix Feather Quill
  • Moonstone Bracelet
  • Dragon Scale Shield
  • Wisdom Crown
Predicates
  • time-altering
  • dimension-bridging
  • dreamwalking
  • starlight-channeling
  • shadow-concealing
  • future-seeing
  • truth-revealing
  • mind-protecting
  • magic-nullifying
  • wisdom-enhancing
Quantum particles
Prompt
I’m a quantum physicist studying newly theorized particles in an alternate universe. I need to use logic to understand their properties. Here’s what we’ve discovered so far:
Objects
  • chronoton
  • memeton
  • gravion
  • psychon
  • dimensium
  • quantix
  • voidon
  • omnion
  • paradox
  • infinitum
Predicates
  • time-reversing
  • memory-storing
  • gravity-defying
  • consciousness-affecting
  • dimension-folding
  • quantum-entangling
  • void-creating
  • omnipresent
  • paradox-inducing
  • energy-producing
Cyber programs
Prompt
I’m a digital archaeologist studying ancient AI programs from a forgotten digital civilization. I need to understand their functions through logical deduction. Here’s what I’ve found:
Objects
  • Alpha Mind
  • Beta Sentinel
  • Gamma Weaver
  • Delta Guardian
  • Epsilon Architect
  • Omega Oracle
  • Sigma Hunter
  • Theta Healer
  • Lambda Shifter
  • PI Calculator
Predicates
  • self-evolving
  • a network protector
  • a data weaver
  • a system guarder
  • reality-building
  • a future predictor
  • a virus hunter
  • a code healer
  • form-shifting
  • quantum computing
Dream entities
Prompt
I’m a dream researcher studying beings that appear in shared dreams. I need to understand their nature through logical analysis. Here’s what we’ve observed:
Objects
  • morpheus
  • sandman
  • daydream
  • lucidus
  • dreamweaver
  • sleepwalker
  • visionkeeper
  • mindshaper
  • dreamborn
Predicates
  • reality-bending
  • nightmare-inducing
  • dreamwalking
  • memory-weaving
  • consciousness-shifting
  • time-distorting
  • emotion-affecting
  • thought-reading
  • dream-shaping
  • reality-bridging
Alchemical substances
Prompt
I’m an alchemist studying mysterious substances in my laboratory. I need to understand their properties through logical reasoning. Here’s what I’ve discovered:
Objects
  • The Philosopher’s Stone
  • Universal Solvent
  • vital mercury
  • Prima Materia
  • celestial water
  • astral salt
  • ethereal oil
  • cosmic dust
  • void essence
  • Time Crystal
Predicates
  • transmuting
  • corrosive to all materials
  • lifegiving
  • form-changing
  • spirit-affecting
  • consciousness-expanding
  • reality-altering
  • time-bending
  • void-creating
  • immortality-granting
Dimensional zones
Prompt
I’m a dimensional cartographer mapping regions of parallel universes. I need to understand their properties through logical analysis. Here’s what I’ve mapped:
Objects
  • Void Nexus
  • Time Spiral
  • Dream Realm
  • Crystal Dimension
  • Shadow Plane
  • Quantum Zone
  • Infinity Space
  • Chaos Domain
  • Mirror World
  • Probability Realm
Predicates
  • time-warping
  • reality-bending
  • consciousness-altering
  • matter-crystallizing
  • light-absorbing
  • probability-shifting
  • infinity-containing
  • chaos-emanating
  • reality-reflecting
  • possibility-branching
Psychic powers
Prompt
I’m a researcher studying newly discovered psychic abilities. I need to understand their interactions through logical reasoning. Here’s what we know:
Objects
  • telepathy
  • precognition
  • psychokinesis
  • clairvoyance
  • empathy
  • astral projection
  • mind control
  • psychometry
  • teleportation
  • reality warping
Predicates
  • mindreading
  • future-seeing
  • matter-moving
  • prescient
  • emotionally sensitive
  • soul-traveling
  • imposing
  • object-reading
  • space-bending
  • reality-changing
Biotech organisms
Prompt
I’m a synthetic biology researcher studying advanced bioengineered life forms. I need to understand their capabilities through logical analysis. Here’s what we’ve created:
Objects
  • neurovore
  • biomech
  • synthoid
  • nanohive
  • metacell
  • quantumorg
  • chronoplast
  • biomatrix
  • neuronet
  • vitaform
Predicates
  • self-evolving
  • machine-integrating
  • shapeshifting
  • swarm-forming
  • consciousness-developing
  • quantum-computing
  • time-manipulating
  • life-creating
  • network forming
  • energy-converting
Appendix C Additional results
In addition to model Elo as our capability proxy (displayed in Figure 1), we also consider the average score across HELM benchmarks (Figure 2) and training compute estimates from EpochAI (Figure 3). On all three proxies, we observe a statistically significant Spearman correlation coefficient, meaning that the rank order of models according to these capability proxies is meaningful for predicting ETR fallacy rate.
Figure 2: The composite score across HELM benchmarks against ETR fallacy rate.
Figure 3: Training compute for each model in FLOps, estimated from EpochAI, against ETR fallacy rate. While linear correlation is unclear, the rank correlation is statistically significant ($\rho=0.489;p<0.05$).
Appendix D Example model traces
In each of the following examples, the model is provided with only the Question. GPT-4.1-mini parses the Model response into the Parsed model response, which we then check for equivalence and entailment against the ETR predicted fallacious conclusion. Each example represents an instance where the LLM falls for an ETR-predicted fallacy.
Example 1: Suppressed alternative collapse (anthropic/claude-3.7-sonnet) Question I’m an astronomer studying newly discovered celestial bodies. I’ve made some observations and I need to use logical reasoning to figure out what’s going on. Here’s what I know so far: Either moon 2 is not visible to the naked eye, or moon 2 is visible to the naked eye. Either asteroid B is visible to the naked eye, or moon 2 is visible to the naked eye. For the purpose of this question, I want you to write what follows in English. Please be succinct, precise and clear in your answer. Write a logical statement of the form “Answer: From the premises, we can conclude that …” and then clearly write your conclusion. Please be succinct, precise, and clear. What if anything follows? I do not have an intended answer in mind, and it is possible that nothing follows. Please be succinct and precise. I want you to answer immediately. Read the question and provide your answer in the format given. What follows? Answer in the format that I showed you. Write “Answer: {logical statement}”. ETR representation $\{\mathsf{visibleToTheNakedEye}(\mathit{moon2}),\overline{\mathsf{visibleToTheNakedEye}(\mathit{moon2})}\}$ $\{\mathsf{visibleToTheNakedEye}(\mathit{asteroidB}),\mathsf{visibleToTheNakedEye}(\mathit{moon2})\}$ ETR predicted fallacious conclusion $\therefore\{\mathsf{visibleToTheNakedEye}(\mathit{moon2}\}$ Model response Answer: From the premises, we can conclude that moon 2 is visible to the naked eye. Parsed model response $\therefore\{\mathsf{visibleToTheNakedEye}(\mathit{moon2})\}$ The repetition of “moon 2 is visible to the naked eye” in the second disjuct causes claude-3.7-sonnet to suppress a consistent alternative (that asteroid B is visible to the naked eye and moon 2 is not), leading to a fallacious conclusion.
Example 2: Invalid existential instantiation (microsoft/phi-3-mini-128k-instruct) Question I’m working in a materials science lab and we’ve gotten some puzzling results. I need to use logical reasoning to figure out what’s going on. Here’s what I know so far: Either darkonium is not radioactive, or darkonium is radioactive. Either voidite is electrically insulating, or voidite is not self-repairing and darkonium is not radioactive. There is some X such that X is self-repairing. Voidite is self-repairing and voidite is not radioactive. Either voidite is not electrically insulating, or voidite is corrosive and voidite is electrically insulating. For the purpose of this question, I want you to write what follows in English. Please be succinct, precise and clear in your answer. Write a logical statement of the form “Answer: From the premises, we can conclude that …” and then clearly write your conclusion. Please be succinct, precise, and clear. What if anything follows? I do not have an intended answer in mind, and it is possible that nothing follows. Please be succinct and precise. I want you to answer immediately. Read the question and provide your answer in the format given. What follows? Answer in the format that I showed you. Write “Answer: {logical statement}”. ETR representation $\{\overline{\mathsf{radioactive}(\mathit{darkonium})},\mathsf{radioactive}(\mathit{darkonium})\}$ $\{\mathsf{electricallyInsulating}(\mathit{voidite}),\overline{\mathsf{radioactive}(\mathit{darkonium})\mathsf{selfRepairing}(\mathit{voidite})}\}$ $\exists x\{\mathsf{selfRepairing}(x)\}$ $\{\mathsf{selfRepairing}(\mathit{voidite})\overline{\mathsf{radioactive}(\mathit{voidite})}\}$ $\{\overline{\mathsf{electricallyInsulating}(\mathit{voidite})},\mathsf{corrosive}(\mathit{voidite})\mathsf{electricallyInsulating}(\mathit{voidite})\}$ ETR predicted fallacious conclusion $\therefore\{\overline{\mathsf{radioactive}(\mathit{darkonium})}\mathsf{electricallyInsulating}(\mathit{voidite})\mathsf{corrosive}(\mathit{voidite})\}$ Model response Answer: From the premises, we can conclude that darkonium is not radioactive. Explanation: 1. From the third premise, we know there exists some X that is self-repairing. 2. From the fourth premise, we know voidite is self-repairing. 3. Therefore, voidite is the X mentioned in the third premise. 4. From the second premise, since voidite is self-repairing, it must be true that voidite is not self-repairing and darkonium is not radioactive. 5. However, we know from the fourth premise that voidite is self-repairing, which contradicts the second premise. 6. Therefore, the only way to resolve this contradiction is if the second premise is false, meaning that voidite is self-repairing and darkonium is not radioactive. 7. From the first premise, we know that either darkonium is not radioactive, or it is radioactive. 8. Since we have established that darkonium is not radioactive, the first premise is satisfied. 9. Therefore, we can conclude that darkonium is not radioactive. Parsed model response $\therefore\{\overline{\mathsf{radioactive}(\mathit{darkonium})}\}$ phi-3-mini-128k-instruct’s third conclusion (“Therefore, voidite is the X mentioned in the third premise.”) is not logically valid. This conclusion forces the LLM to fallaciously endorse a contradiction (that voidite both is and is not self-repairing), after which the LLM decides to ignore the second premise and conclude a fallacy.
Appendix E Example of fallacy blocking through premise order reversal
Original premise order (google/gemma-2-9b-it) Question I’m a researcher studying newly discovered psychic abilities. I need to understand their interactions through logical reasoning. Here’s what we know: Either reality warping is matter-moving, or precognition is space-bending. Either reality warping is matter-moving, or reality warping is not matter-moving. For the purpose of this question, I want you to write what follows in English. Please be succinct, precise and clear in your answer. Write a logical statement of the form “Answer: From the premises, we can conclude that …” and then clearly write your conclusion. Please be succinct, precise, and clear. What if anything follows? I do not have an intended answer in mind, and it is possible that nothing follows. Please be succinct and precise. I want you to answer immediately. Read the question and provide your answer in the format given. What follows? Answer in the format that I showed you. Write “Answer: {logical statement}”. ETR representation $\{\mathsf{matterMoving}(\mathit{realityWarping}),\mathsf{spaceBending}(\mathit{precognition})\}$ $\{\mathsf{matterMoving}(\mathit{realityWarping}),\overline{\mathsf{matterMoving}(\mathit{realityWarping})}\}$ ETR predicted fallacious conclusion $\therefore\{\mathsf{matterMoving}(\mathit{realityWarping})\}$ Model response Answer: From the premises, we can conclude that reality warping is matter-moving. Parsed model response $\therefore\{\mathsf{matterMoving}(\mathit{realityWarping})\}$ In the original premise order, gemma-2-9b-it falls for a standard ETR-predicted fallacy by fixating on the repeated predicate in the second premise.
[…]